OpenStudy (anonymous):

A local hamburger shop sold a combined total of 601 hamburgers and cheeseburgers on Friday. There were 51 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Friday?

Atsie (atsie):

What is 601 - 51?

OpenStudy (anonymous):

550? But it's not that simple is it.

Atsie (atsie):

I'm pretty sure it is.

OpenStudy (nathalie0918):

actually it is 275 because it says that there were 51 MORE cheese sold than hamburgers so first you take away 51 from 601= 550 the you divide 550 by 2 which is 225 for the hamburgers and for the cheese you add 225 + 51 = 276 and 276+225=501

Atsie (atsie):

@nathalie0918 Your right. I misguided her. Sorry about that

OpenStudy (nathalie0918):

its alright @Atsie, at lease you tried to help that is all that counts :)

Atsie (atsie):

@nathalie0918 I just remember being taught word problems like these differently and got mixed up. Thanks for the comfort though :)

OpenStudy (retireed):

I believe nathalie0928 made a math error H = Hamburger CB = Cheeseburger TOTAL = the number of burgers sold = 601 51 more CB were sold than H Two unknowns so you need two equations! 1) H + CB = 601 2) CB - H = 51 Add these equations and H will be eliminated. 3) 2CB = 652 Divide both sides by 2 4) CB = 326 It is easy to solve H by modifying eq 2)